### 01 Mar 2024 v1

# End-to-end Graph-Sequential Representation Learning for Accurate Recommendations

Abstract.
Many recent advancements in recommender systems have focused on developing sequence-based and graph-based approaches. Both approaches proved useful in modeling intricate relationships within behavioral data, leading to promising outcomes in personalized ranking and next-item recommendation tasks while maintaining good scalability. However, they capture very different signals from data. While the former approach represents users directly through ordered interactions with recent items, the latter one aims to capture indirect dependencies across the interactions graph. This paper presents a novel multi-representational learning framework that exploits the synergies between these two paradigms. Our empirical evaluation on several datasets demonstrates that mutual training of sequential and graph components with the proposed framework significantly improves recommendations performance.
Recommender Systems, Sequential Learning, Graph Neural Networks, Contrastive Learning
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: The ACM Web Conference 2024; May 13–17, 2024; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
Figure 1. The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encode input item and user IDs. (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations. (v) Fusing Layer capable of mixing representations from both encoders. (vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) contrastive objective used for model training.
MRGSRec model overview.
Figure 1. The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encode input item and user IDs. (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations. (v) Fusing Layer capable of mixing representations from both encoders. (vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) contrastive objective used for model training.
1. Introduction
The primary objective of recommender systems is to offer users personalized recommendations aligned with their interests by leveraging past engagements between users and items. In recent years, deep learning-based solutions have emerged as dominant players in the evolution of recommender systems, causing the rapid development of diverse techniques and methods (Zhang et al., 2019; Gao et al., 2023). These solutions can construct accurate and expressive representations of users and items and address various recommendations tasks (Kang and McAuley, 2018; Sun et al., 2019; He et al., 2020). In most cases, input data for such models may be represented as a user-specific sequence of recently interacted items. Leveraging only this information, however, has its limitations: concerned with individual user interactions, this approach captures patterns only for direct dependencies, which may lack sensitivity to a broader context provided by additional interdependencies across entities. On the contrary, Graph Neural Networks (GNNs) (Wang et al., 2019; He et al., 2020) networks provide a powerful way of explicitly modeling information through indirect dependencies, but adapting them to sequential setup can be cumbersome. Each approach therefore addresses the same problem of learning user preferences by leveraging patterns from independent data representations. To our knowledge, only one study has attempted to integrate both representations within a single model (Wang et al., 2022). However, it was constrained by a limited encoding of sequential interactions. In contrast, we aim to design a general approach, allowing for the use of various versions of sequential and graph-based components, constrained only by encoder signatures. Our contributions are as follows:
  • We propose MRGSRec (Multi-Representation Graph-based and Sequence-based approach for Recommendations), novel architecture design that integrates graph-based and sequential representations into an end-to-end framework;
  • We present a scheme for knowledge exchange between different representations based on contrastive learning;
  • We assemble a composite loss function that maximizes the cumulative effect of all components of the architecture.
2. Related work
We categorize existing models into three types: sequence-based, graph-based, and multi-representational, integrating both categories into one algorithm.
Sequential Recommendations
Transformer architectures (Vaswani et al., 2023) have proven to be effective across many domains (Brown et al., 2020; Dosovitskiy et al., 2021; Carion et al., 2020), including their utility in sequence modeling. Their unique properties have contributed to their success in various applications, including recommender systems, leading to the development of top-performing methods. Examples include SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) demonstrating state-of-the-art performance. The effectiveness of these attention-based architectures in modeling sequential data sparked many further improvements and adaptations aimed at enhancing recommendations quality (Xie et al., 2021; Qiu et al., 2022; Klenitskiy and Vasilev, 2023). Notably, the usage of cross-entropy loss for SASRec architecture (SASRec+) (Klenitskiy and Vasilev, 2023) has demonstrated superior performance compared to other techniques.
Graph-based Recommendations
An alternative approach to match user preferences involves using graph representations from user-item interactions. Graph Convolution Networks (GCNs), notably NGCF (Wang et al., 2019) and LightGCN (He et al., 2020), have been recognized as highly effective solutions within this domain, with NGCF being foundational and LightGCN serving as both an ancestor and an enhancement.
Multi-Representational Recommendations
Both previous ideas have their limitations: sequential recommendations excel at extracting patterns from user-specific interactions but overlook broader interactions. Conversely, GCNs grasp the indirect interdependencies but struggle to leverage them within user-specific history as precise as sequential models. To the best of our knowledge, only one approach, namely MCLSR (Wang et al., 2022) attempts to bridge this gap by using both paradigms, however, only during the training process, by propagating patterns from indirect interdependencies into item embedding representations. Our approach, MRGSRec, enhances the idea from MCLSR of fusing signals across different representations, by utilizing both sequence-based and graph-based representations during the whole process of training and inference. This design offers an end-to-end framework for combining various architectures.
3. Proposed approach
This section presents the problem statement for the considered recommendation scenario and explains in detail the components of the MRGSRec framework, providing insights into their roles in enhancing model quality. Model overview is shown in Fig. 1.
We consider a standard next-item recommendation task. The goal is to establish relations between a set $\mathcal{U}$ of $M$ users and a set $\mathcal{I}$ of $N$ items in a sequential manner, i.e., for a given user $u\in\mathcal{U}$ described by a sequence of the previously interacted items $S_{u}=(i^{u}_{1},\,i^{u}_{2},\,\dots,\,i^{u}_{|S_{u}|})$, $i^{u}_{k}\in\mathcal{I}$, the task is to predict the next item $i^{u}_{|S_{u}|+1}$ the user is likely to interact with. Below we outline the components of our solution.
3.1. Embedding layers
Given a user ID $u$ and an interacted sequence $S_{u}$ of item IDs, we keep only the last $c$ interactions from $S_{u}$. We then convert the user and remaining item IDs into embeddings using learnable embedding layers $U_{e}:\mathbb{N}\rightarrow\mathbb{R}^{d}$ and $I_{e}:\mathbb{N}\rightarrow\mathbb{R}^{d}$ for user ID and item IDs respectively. This process produces a user embedding $e^{u}=U_{e}(u)\in\mathbb{R}^{d}$ and a sequence of item embeddings $\left(e^{u}_{1},e^{u}_{2},\dots,e^{u}_{c}\right)=E^{u}\in\mathbb{R}^{c\times d}$, where $e^{u}_{k}=I_{e}(i^{u}_{k})$ and $d$ denotes the embedding size. To enhance item representation with position-specific information, we add learnable positional embeddings similarly to existing approaches (Kang and McAuley, 2018; Sun et al., 2019; Xie et al., 2021; Qiu et al., 2022).
3.2. Sequential Encoder
To capture the direct interactions, we utilize a sequential encoder block within the MRGSRec framework. This block is tasked with producing two representations: an aggregated local behavioral representation and local enriched embeddings for all input items, with dimensions $\mathbb{R}^{d}$ and $\mathbb{R}^{c\times d}$ respectively. Without integrating user-specific data, the encoder output lacks personalized context, producing user-agnostic output. We address this limitation by concatenating $e^{u}$ and $E^{u}$ before applying the encoder, allowing the sequential encoder to enrich each item representation with user-specific context. We then extract the local behavioral representation $e_{l}^{u}$ and local enriched embeddings $E_{l}^{u}$ from the encoder output as specified in the following equation:
(1) $e_{l}^{u}\,\|\,E_{l}^{u}=\text{SequentialEncoder}\left(e^{u}\,\|\,E^{u}\right),$
where SequentialEncoder is an encoder part from the transformer architecture (Vaswani et al., 2023), that processes an input matrix of size $(c+1)\times d$ to produce an enriched representation of the same size.
3.3. Graph Encoder
To incorporate the ability to capture indirect interdependencies, we propose employing GCNs, a method used in various prior works (Gao et al., 2023; Wang et al., 2019; He et al., 2020). Within the MRGSRec framework, applying a graph encoder requires similar constraints as discussed in the sequential encoder section. Using GCNs involves introducing two input matrices: the initial representation matrix for all users and items $E^{(0)}_{g}\in\mathbb{R}^{(M+N)\times d}$ and the graph adjacency matrix $A\in\mathbb{R}^{(M+N)\times(M+N)}$. To construct $E^{(0)}_{g}$, we utilize embeddings for all users and items from the previously defined embedding layers $U_{e}$ and $I_{e}$ by concatenating them. For adjacency matrix initialization, we follow the approach from prior works (Wang et al., 2019; He et al., 2020). Initially, we introduce the interaction matrix $R\in\mathbb{R}^{M\times N}$, in which elements $r_{ij}$ are set to 1 when there is an interaction between user $i$ and item $j$, and 0 otherwise. Using matrix $R$, we can define matrix $A$ in the block notation:
(2) $A=\begin{pmatrix}0&R\\ R^{\top}&0\\ \end{pmatrix}.$
To apply graph encoder, we use $E^{(0)}_{g}$ and $A$ in the recursive equation:
(3) $E^{(k)}_{g}=\text{GraphEncoder}\left(E_{g},\,A\right)=D^{-\frac{1}{2}}\,A\,D^{-\frac{1}{2}}\,E^{(k-1)}_{g}\,,$
where $A$ denotes the graph’s adjacency matrix, $D_{ii}=\sum_{j=1}^{M+N}{a_{ij}}$ stands for the diagonal degree matrix of $A$, $k$ represents the number of graph encoder layers, and $E^{(k)}_{g}$ signifies the output of the graph encoder, which provides enriched representations for all users and items. To obtain the global behavioral representation $e^{u}_{g}\in\mathbb{R}^{d\times 1}$ and global enriched embeddings $E^{u}_{g}\in\mathbb{R}^{c\times d}$ following the graph encoder’s execution, we derive the final representation from $E^{(k)}_{g}$ using the IDs of given user $u$ and each item in $S_{u}$.
3.4. Fusion Layer
After processing input data with both sequential and graph encoders, we obtain pairs of embeddings $e^{u}_{l}$ and $e^{u}_{g}$ for local and global behavioral representations, respectively. To integrate information from both encoders into our final representation, we utilize a shallow feed-forward block as described in Eq. 4:
(4) $e^{u}_{f}=W_{2}\,\text{ReLU}\left(W_{1}(e^{u}_{l}\|e^{u}_{g})\right),$
where $e^{u}_{f}\in\mathbb{R}^{d}$ denotes the fused behavioral representation. $W_{1}\in\mathbb{R}^{4d\times 2d}$ and $W_{2}\in\mathbb{R}^{d\times 4d}$ are learnable projection matrices, and $\|:\mathbb{R}^{d},\,\mathbb{R}^{d}\rightarrow\mathbb{R}^{2d}$ represents the concatenation function.
3.5. Loss Functions
The essential component of MRGSRec, enabling end-to-end usage of both contexts, lies in its selection of loss functions. Inspired by the cross-domain recommender systems (Xu et al., 2023), we construct objectives for every intermediate step within our framework. For context-specific behavior representations, we use local and global objectives $\mathcal{L}_{l}$ and $\mathcal{L}_{g}$ which are the same as the downstream loss functions from the chosen encoder architecture (e.g., next item prediction from SASRec (Kang and McAuley, 2018) for sequential encoder and BPR-loss plus regularization loss from LightGCN (He et al., 2020) for graph encoder). Additionally, we introduce an auxiliary objective $\mathcal{L}_{f}$ based on fused behavioral representation:
(5) $\mathcal{L}_{f}=\sum_{u\in\mathcal{U}}{-\log\frac{\exp^{(e^{u}_{f})^{\top}I_{e}(i^{u}_{p})}}{\sum_{i_{n}\in\text{Sample}(u)}{\exp^{(e^{u}_{f})^{\top}I_{e}(i_{n})}}+\exp^{(e^{u}_{f})^{\top}I_{e}(i^{u}_{p})}}},$
where $i^{u}_{p}$ is the next interacted item for user $u$, and $i_{n}$ is a randomly selected negative item ID outside of $S_{u}$. Also, this representation is used for computing item relevance scores by using the dot-product between this embedding and item embeddings taken from $I_{e}$.
Finally, we introduce contrastive loss $\mathcal{L}_{c}$ between local and global enriched item representations $E^{u}_{l}$ and $E^{u}_{g}$, aiding the model to make a distinction between direct and indirect dependencies:
(6) $\mathcal{L}_{c}=\sum_{u\in\mathcal{U}}{\sum_{i=1}^{c}{-\log\frac{\exp^{(E^{u}_{l})^{\top}_{i}(E^{u}_{g})_{i}}}{\sum\limits_{j=1}^{c}{\exp^{(E^{u}_{l})^{\top}_{i}(E^{u}_{g})^{\top}_{j}}}}}},$
where $(E^{u}_{g})_{i}$ and $(E^{u}_{l})_{i}$ – global and local enriched representations of $i$-th item in the sequence respectively.
The overall objective is given as follows:
(7) $\mathcal{L}(\theta)=\alpha\mathcal{L}_{l}+\beta\mathcal{L}_{g}+\gamma\mathcal{L}_{f}+\delta\mathcal{L}_{c},$
where $\alpha,\,\beta,\,\gamma,\,\delta$ are model hyper-parameters determining each objective’s influence on the overall model objective.
4. Experimental setup
Datasets
We experiment on four benchmark datasets: Amazon-Beauty, Amazon-Clothing, Amazon-Sports, and MovieLens-1M. Data preprocessing follows standard procedure: treating all interactions as positives, filtering out users and items with fewer than 5 interactions, and sorting user sequences by interactions timestamp. Table 1 summarizes statistics after preprocessing.
Dataset #Items #Users #Interactions Avg. length
Beauty 12,101 22,363 198,502 8.876
Clothing 23,033 39,387 278,677 7.075
Sports 18,357 35,598 296,337 8.325
ML-1M 3,706 6,040 1,000,209 165.598
Table 1. Statistics of the experimental datasets
Evaluation setting
[…]
Dataset Metric SASRec BERT4Rec SASRec+ NGCF LightGCN MCLSR MRGSRec Uplift (%)
Beauty HR@5 0.0352 0.0182 0.0426 0.0130 0.0154 0.0371 0.0441 3.52
HR@10 0.0611 0.0412 0.0663 0.0215 0.0253 0.0614 0.0681 2.71
NDCG@5 0.0214 0.0183 0.0289 0.0070 0.0079 0.0212 0.0301 4.15
NDCG@10 0.0273 0.0254 0.0343 0.0094 0.0098 0.0275 0.0349 1.75
Clothing HR@5 0.0159 0.0137 0.0207 0.0077 0.0089 0.0168 0.0221 6.76
HR@10 0.0271 0.0206 0.0339 0.0133 0.0156 0.0287 0.0352 3.83
NDCG@5 0.0090 0.0078 0.0114 0.0041 0.0043 0.0093 0.0121 6.15
NDCG@10 0.0116 0.0105 0.0155 0.0059 0.0071 0.0117 0.0162 4.52
Sports HR@5 0.0216 0.0178 0.0275 0.0115 0.0133 0.0233 0.0289 5.09
HR@10 0.0324 0.0322 0.0417 0.0155 0.0179 0.0318 0.0434 4.08
NDCG@5 0.0135 0.0102 0.0176 0.0068 0.0072 0.0134 0.0186 5.68
NDCG@10 0.0167 0.0153 0.0218 0.0073 0.0080 0.0168 0.0227 4.12
ML-1M HR@5 0.1067 0.0745 0.1244 0.0439 0.0485 0.1077 0.1312 5.47
HR@10 0.1804 0.1333 0.2133 0.0864 0.0921 0.1813 0.2248 5.39
NDCG@5 0.0639 0.0439 0.0761 0.0286 0.0307 0.0634 0.0829 8.94
NDCG@10 0.0918 0.0623 0.1065 0.0477 0.0549 0.0921 0.1134 6.48
Table 2. Experimental results on benchmark datasets.
Baseline Algorithms
We compare MRGSRec with baseline algorithms from both sequence-based and graph-based approaches. Baseline models include SASRec, BERT4Rec, and SASRec+ for sequential approaches, and NGCF and LightGCN for graph-based models. Additionally, we evaluate MRGSRec against MCLSR, an approach integrating both data representations. During this comparison, MRGSRec implementation incorporates the sequential encoder and loss function from SASRec+, and the graph encoder with objective function from LightGCN to explore its boost upon these strong baselines. All information about hyper-parameter settings can be found in our GitHub repository  [^1] https://github.com/NonameUntitled/MRGSRec.
5. Results and Discussion
Table 2 reports the overall performance compared with baselines. The obtained results highlight a more pronounced relative uplift for $n=5$ compared to $n=10$. Although there is also an uplift for $n=20$ (not reported here for brevity), it is observed to be of a lesser extent. We conclude that the model excels at promoting positive items to the top of recommendation lists, but revealing additional positive examples undetected by other models remains challenging.
Our analysis shows that the ML-1M dataset, with a higher density of interactions, demonstrates the most significant relative improvement in the NDCG metric. This suggests that our model may enhance performance when dealing with more interactions, possibly due to graph models performing better for frequently consumed items. This observation could explain the comparatively lower uplift metric values observed for the Beauty dataset.
In this paper, we have presented a novel end-to-end framework for leveraging sequence-based and graph-based data representations into a single model to improve the representation of user behaviors. The proposed framework, MRGSRec, constructs two independent representations from sequence-based and graph-based encoders that can capture different patterns in the input data. Our experiments on classical benchmarks demonstrate the effectiveness of our approach against diverse families of state-of-the-art methods.

### 14 Mar 2024 v2

@@
 Abstract.
-Many recent advancements in recommender systems have focused on developing sequence-based and graph-based approaches.
+Recent recommender system advancements have focused on developing sequence-based and graph-based approaches.
 Both approaches proved useful in modeling intricate relationships within behavioral data, leading to promising outcomes in personalized ranking and next-item recommendation tasks while maintaining good scalability.
 However, they capture very different signals from data.
-While the former approach represents users directly through ordered interactions with recent items, the latter one aims to capture indirect dependencies across the interactions graph.
-This paper presents a novel multi-representational learning framework that exploits the synergies between these two paradigms.
+While the former approach represents users directly through ordered interactions with recent items, the latter aims to capture indirect dependencies across the interactions graph.
+This paper presents a novel multi-representational learning framework exploiting these two paradigms’ synergies.
 Our empirical evaluation on several datasets demonstrates that mutual training of sequential and graph components with the proposed framework significantly improves recommendations performance.
 Recommender Systems, Sequential Learning, Graph Neural Networks, Contrastive Learning
-copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: The ACM Web Conference 2024; May 13–17, 2024; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
+journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3651499isbn: 979-8-4007-0172-6/24/05ccs: Information systems Recommender systems
 Figure 1.
-The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encode input item and user IDs.
+The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encodes input item and user IDs.
 (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations.
 (v) Fusing Layer capable of mixing representations from both encoders.
-(vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) contrastive objective used for model training.
-MRGSRec model overview.
+(vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) Contrastive Objective used for model training
+MRGSRec model overview
 Figure 1.
-The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encode input item and user IDs.
+The proposed MRGSRec model consists of nine blocks, colored in yellow and red: (i) Item Embedding Layer and (ii) User Embedding Layer that encodes input item and user IDs.
 (iii) Sequential Encoder and (iv) Graph Encoder that aggregate information leveraging different data representations.
 (v) Fusing Layer capable of mixing representations from both encoders.
-(vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) contrastive objective used for model training.
+(vi) Local Objective, (vii) Global Objective, (viii) Fusing Objective, and (ix) Contrastive Objective used for model training
 1. Introduction
@@
 Leveraging only this information, however, has its limitations: concerned with individual user interactions, this approach captures patterns only for direct dependencies, which may lack sensitivity to a broader context provided by additional interdependencies across entities.
-On the contrary, Graph Neural Networks (GNNs) (Wang et al., 2019; He et al., 2020) networks provide a powerful way of explicitly modeling information through indirect dependencies, but adapting them to sequential setup can be cumbersome.
+On the contrary, Graph Neural Networks (GNNs) (Wang et al., 2019; He et al., 2020) provide a powerful way of explicitly modeling information through indirect dependencies, but adapting them to sequential setup can be cumbersome.
 Each approach therefore addresses the same problem of learning user preferences by leveraging patterns from independent data representations.
 To our knowledge, only one study has attempted to integrate both representations within a single model (Wang et al., 2022).
-However, it was constrained by a limited encoding of sequential interactions.
-In contrast, we aim to design a general approach, allowing for the use of various versions of sequential and graph-based components, constrained only by encoder signatures.
+However, it was constrained by a sequential encoder architecture.
+In contrast, we aim to design a general approach, allowing the useage of various versions of sequential and graph-based components, constrained only by encoder signatures.
 Our contributions are as follows:
@@
 2. Related work
-We categorize existing models into three types: sequence-based, graph-based, and multi-representational, integrating both categories into one algorithm.
+We categorize existing models into three types: sequence-based, graph-based, and multi-representational (models that integrate both categories into one algorithm).
 Sequential Recommendations
@@
 Conversely, GCNs grasp the indirect interdependencies but struggle to leverage them within user-specific history as precise as sequential models.
-To the best of our knowledge, only one approach, namely MCLSR (Wang et al., 2022) attempts to bridge this gap by using both paradigms, however, only during the training process, by propagating patterns from indirect interdependencies into item embedding representations.
-Our approach, MRGSRec, enhances the idea from MCLSR of fusing signals across different representations, by utilizing both sequence-based and graph-based representations during the whole process of training and inference.
-This design offers an end-to-end framework for combining various architectures.
+To the best of our knowledge, only one approach, namely MCLSR (Wang et al., 2022) attempts to bridge this gap by using both paradigms, however, the fusing between different representations is used only during the training process by propagating patterns from indirect interdependencies into item representations, while during the inference only sequential part is leveraged.
+Also, MCLSR uses an architecture with limited capacity for the sequential part of the model.
+MRGSRec enhances the idea from MCLSR by utilizing both sequence-based and graph-based representations during the whole training and inference process and providing flexible choices of different encoders for both sequence-based and graph-based parts by having only signature constraints.
+By combining various architectures, this design offers an end-to-end solution.
 3. Proposed approach
@@
 1.
-We consider a standard next-item recommendation task.
+We consider a standard next-item prediction task.
 The goal is to establish relations between a set $\mathcal{U}$ of $M$ users and a set $\mathcal{I}$ of $N$ items in a sequential manner, i.e., for a given user $u\in\mathcal{U}$ described by a sequence of the previously interacted items $S_{u}=(i^{u}_{1},\,i^{u}_{2},\,\dots,\,i^{u}_{|S_{u}|})$, $i^{u}_{k}\in\mathcal{I}$, the task is to predict the next item $i^{u}_{|S_{u}|+1}$ the user is likely to interact with.
@@
 This process produces a user embedding $e^{u}=U_{e}(u)\in\mathbb{R}^{d}$ and a sequence of item embeddings $\left(e^{u}_{1},e^{u}_{2},\dots,e^{u}_{c}\right)=E^{u}\in\mathbb{R}^{c\times d}$, where $e^{u}_{k}=I_{e}(i^{u}_{k})$ and $d$ denotes the embedding size.
-To enhance item representation with position-specific information, we add learnable positional embeddings similarly to existing approaches (Kang and McAuley, 2018; Sun et al., 2019; Xie et al., 2021; Qiu et al., 2022).
+To enhance item representation with position-specific information, we add learnable positional embeddings similar to existing approaches (Kang and McAuley, 2018; Sun et al., 2019; Xie et al., 2021; Qiu et al., 2022).
 3.2. Sequential Encoder
@@
 We address this limitation by concatenating $e^{u}$ and $E^{u}$ before applying the encoder, allowing the sequential encoder to enrich each item representation with user-specific context.
-We then extract the local behavioral representation $e_{l}^{u}$ and local enriched embeddings $E_{l}^{u}$ from the encoder output as specified in the following equation:
+As specified in the equation above, then we extract local behavioral representation $e_{l}^{u}$ and local enriched embeddings $E_{l}^{u}$:
 (1) $e_{l}^{u}\,\|\,E_{l}^{u}=\text{SequentialEncoder}\left(e^{u}\,\|\,E^{u}\right),$
@@
 3.5. Loss Functions
-The essential component of MRGSRec, enabling end-to-end usage of both contexts, lies in its selection of loss functions.
+The essential component of MRGSRec, enabling end-to-end usage of both contexts, lies in its loss functions.
 Inspired by the cross-domain recommender systems (Xu et al., 2023), we construct objectives for every intermediate step within our framework.
@@
 Table 1 summarizes statistics after preprocessing.
+Table 1. Statistics of the experimental datasets
 Dataset #Items #Users #Interactions Avg. length
@@
 ML-1M 3,706 6,040 1,000,209 165.598
-Table 1. Statistics of the experimental datasets
 Evaluation setting
 […]
+Table 2. Experimental results on benchmark datasets
 Dataset Metric SASRec BERT4Rec SASRec+ NGCF LightGCN MCLSR MRGSRec Uplift (%)
@@
 NDCG@10 0.0918 0.0623 0.1065 0.0477 0.0549 0.0921 0.1134 6.48
-Table 2. Experimental results on benchmark datasets.
 Baseline Algorithms
-We compare MRGSRec with baseline algorithms from both sequence-based and graph-based approaches.
-Baseline models include SASRec, BERT4Rec, and SASRec+ for sequential approaches, and NGCF and LightGCN for graph-based models.
+We compare MRGSRec with baselines from both sequence-based and graph-based approaches.
+These models include SASRec, BERT4Rec, and SASRec+ for sequence-based, and NGCF with LightGCN for graph-based models.
 Additionally, we evaluate MRGSRec against MCLSR, an approach integrating both data representations.
 During this comparison, MRGSRec implementation incorporates the sequential encoder and loss function from SASRec+, and the graph encoder with objective function from LightGCN to explore its boost upon these strong baselines.
-All information about hyper-parameter settings can be found in our GitHub repository  [^1] https://github.com/NonameUntitled/MRGSRec.
 5. Results and Discussion